The base of a triangel exceeds the height by 9 inches. If the area is 180 squares inches,find the length of the base and height of the triangle
Answer by Boreal(15235)

height =x
base=x+9
x(x+9)/2=180, because A=(1/2)bh
x^2+9x=360
x^2+9x-360=0
(x+24)(x-15)=0
x=-24 and 15, only 15 is sensible
height is 15 inches, base is 24 inches. Half their product is 180 in^2, the area.